

********************************************************************
UARS WEEKLY STATUS REPORT - 11 DEC 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
Announcing the newest member of the SOLSTICE Family - Maggie Boyle -
Brian and Lisa Boyle's new daughter. Congratulations Brian!
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The instrument is now off as scheduled.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 444 CD-ROM requests as of 11-December-1998.
Two orders were received and shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
18 days HAL (4.80) PROD (L2-L3) (89R-80C=9L)
47 days HAL (4.80) PROD (L3) (1874R-1476C=398L)
46 days SOL (4.62) PROD (L2-L3) (2313R-2198C=115L)
25 days SUS (7.00) PROD (L1) (43R-25C=18L)
46 days SUS (7.00) PROD (L3) (2540R-823C-3F=1714L)
0 days CLA (7.00) ENG (L1-L3) (24R-21C=3L)
One month of old accounting data was successfully loaded from the operational
system archive tape onto the development database. This information
will be analyzed to determine correlation in SOLSTICE processing time
degradation.
Preventive Maintenance for CDHF systems and peripherals is scheduled for
Tuesday 12/15/98.
The CDHF continued the configuration phase of the of UCSS. This effort involves
the software installation of UOAS 5.5, UCSS 10.1, MSS 4.1, and NRT 5.4.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 27 November 1998 (Orbit 39418, GMT Day 332) through 04
December 1998 (Orbit 39522, GMT Day 338).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 331/0006 39403 AXIS On
HALOE 334/1254 39456 On
PEM 334/1619 39459 Off
PEM 335/0336 39465 AXIS Off
WINDII 335/2239 39477 Off
PEM 336/0021 39478 On (PMON)
A summary of Instrument operations supplied by the Mission Planning Group
for the month of December 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle ranged between 65.8 degrees to 37.0 degrees for this report
period. The beta angle is now decreasing toward a minimum angle of -1.0
degrees on 13 December 1998 (DOY 347).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 to 1.6 degrees C. Additional battery performance data
for each battery on 27 November 1998 and 04 December 1998 are provided in
Appendix D.
The clock error ranged from 10.4 to -11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 04
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HALOE 03
% PEM 04
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 07
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496917
Transponder B: 98/300 - 2287.498980
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/337 show the observatory data loss to be
21 hours, 40 minutes, 6.772 seconds (an increase of 26.752 seconds since
last report period). This is a 0.0343 percent data loss which equals a
99.9657 percent data capture for the mission.
The increase of 26.752 seconds occurred over five (5) days:
7 5.216 seconds on Day 331 due to hits in data (IP);
7 9.888 seconds on Day 332 due to hits in data (IP);
7 7.904 seconds on Day 333 due to hits in data (IP);
7 3.520 seconds on Day 334 due to hits in data (IP); and
7 0.224 seconds on Day 336 due to data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Preparation of 0.5 degrees (colds side down) roll procedure to be executed
in December 1998 and January 1999 continues.
The procedure for the clock change to include the 01 January 1999 leap
second adjustment continues.
Flight Operations and SOLSTICE have been developing the methodology for
SOLSTICE Day-Only Operation.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - No change.
Flight Dynamics Facility (FDF) - see CMS paragraph.
IP NETWORK TESTING - Complete
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Testing of the Trending System v.1.1 release (CM-0009) continued.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 13
December 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 December 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39403 331/0006 98-473 PEM Power AXIS ON
39408 331/0825 98-474 WINDII Error clear
39416 331/2032 98-475 OBC Clk Adj fr 40.2 to 20.6 msec/day
39435 333/0331 98-476 OBC Clk Adj fr 20.6 to 0.00 msec/day
39452 334/0654 98-477 OBC Clk Adj fr 0.00 to 20.6 msec/day
39456 334/1254 98-478 HALOE Power ON
39459 334/1619 98-479 PEM Power OFF
39465 335/0336 98-480 PEM AXIS Pwr OFF
39469 335/1034 98-481 WINDII Error clear
39477 335/2239 98-482 WINDII Power OFF
39478 336/0021 98-483 PEM PMON GROUP 10 ENA
39490 336/1913 98-484 HALOE MUB DISA (HIGH BARREL TEMP.)
39490 336/1948 98-484 HALOE MUB ENA (BARl. TEMP BACK IN LIM.)
39492 336/2243 98-485 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 December 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP Y-LO HRDI BARREL TEMP 39414-39417
HACS4 Y-HI Control status word 4 39490
HRPTSBARTMP R-LO HRDI BARREL TEMP 39490
HACS4 R-HI Control status word 4 39490
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIERRLIMIT WINDII Software Error Code = E03 39408
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 December 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39435 333/03:36 (29 Nov 98) from 20.6 to 0.00 msec/day
39452 334/07:06 (30 Nov 98) from 0.00 to 20.6 msec/day
39492 336/22:55 (02 Dec 98) from 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 December 1998
27 Nov 1998 - GMT Day 331 Beta = 65.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +95.2/+67.2 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +11.2/-33.6 +3.87/+2.60 +22.8/-15.2 26.7 1.030 17.3
3 +44.8/ 0.0 +2.92/+1.66 +22.0/-14.4 26.7 1.031 16.7
04 Dec 1998 - GMT Day 338 Beta = 37.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +84.0/+56.0 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +28.0/-22.4 +5.47/+4.51 +29.6/-15.2 25.8 1.027 21.0
3 +56.0/-5.6 +4.51/+2.92 +28.8/-15.2 25.8 1.026 20.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 December 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT 39407 331/0642 14170 Generic Late Acquisition
98-092 FOT 39415 331/1858 14170 Generic Late Acquisition
98-092 FOT 39420 332/0327 14170 Generic Late Acquisition
98-115 FOT 39433 333/0009 N/A TAC 4 Crashed
98-116 FOT 39468 335/0852 20415 IP Problem at White Sands
98-116 FOT 39469 335/0937 20415 IP Problem at White Sands
98-119 FOT 39492 336/2243 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT NOV 98 14170 GENERIC LATE ACQ
98-107 FOT 39308 324/16:40 TAC 2 PROB
98-108 FOT 39309 324/17:23 TAC PROBLEMS
98-109 FOT 39323 325/15:46 TAC PROBLEMS
98-110 FOT 39325 325/19:16 TAC PROBLEMS
98-111 FOT 39341 326/20:52 20396 DESKTOP DOWN
98-112 FOT 39342 326/22:04 20396 DESKTOP DOWN
98-113 FOT 39342 326/22:43 CMD PROBLEMS
98-114 FOT 39355 327/17:35 TAC/DMP PROBLEMS
98-115 FOT 38433 333/00:09 TAC 4 CRASH
98-116 FOT 39648 335/08:52 20415 IP DOWN
98-118 FOT 39490 336/19:18 BANK DMP TIME OUT
98-119 FOT 39490 336/19:48 BANK DMP TIME OUT
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-117 FOT DEC 98 See List 14170 GENERIC LATE ACQ
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 04 December 1998
INSTRUMENT SUMMARY FOR MONTH OF December 1998 (hours)
UNPLANNED UNSCHEDULED
ON Day- OFF OFF OFF
Only (Insufficient (Other
Instrument (Actual) Ops (Scheduled) Power) Problem) COMMENTS
---------- ------ --- ----------- ----------- ----------- -------------
ACRIM 720 0 0
HALOE 224 496 0 0
HRDI 247 473 0 0 0
MLS 218 502 0 0
PEM (axis) 88 632 0 0
PEM (zeps) 178 534 8 0 0
PEM (heps) 0 0 0 720 See Note 1
SOLSTICE 720 0 0 0
SUSIM 720 0 0 0
WINDII 360 269 91 0
Note 1. PEM High Energetic Particle System (HEPS) was scheduled to be
ON for day-light only operation. However, HEPS remained OFF due to a
problem encountered with automatic turn-on procedure for instrument high
voltage power.
Cold side of spacecraft was facing northward: 28.5 days
Cold side of spacecraft was facing southward: 1.5 days
SEVNTFW49